Which is larger for a hoop of mass M and radius R, that is rolling without slipping - its translational or rotational kinetic energy? 1.(A) Translational kinetic energy 1.(B) Rotational kinetic energy 1.(C) Both are the same 1.(D) Answer depends upon radius

  1. Which is larger for a hoop of mass M and radius R, that is rolling without slipping - its translational or rotational kinetic energy?
1.(A) Translational kinetic energy
1.(B) Rotational kinetic energy
1.(C) Both are the same
1.(D) Answer depends upon radius

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
 
Ktrans/ Krot = (1/2* m* v^2) / ½ * I(hoop)* w^2 = m * v^2 / m* v^2 = 1
 
Therefore, the translational and rotational kinetic energies are the same.
 
 
Regards
Arun (askIITians forum expert)